You are here: Foswiki>PartnerObserving Web>TechNightLogs>December22UT (21 Jul 2009, NormCushing)Edit Attach

PartnerObserving Log

morning Andrea, Fernando & Vincenzo wrapping up. In the User Interface, Andrea changed the OB Progress to show number of images taken out of the total instead of number of dithers out of total. This was needed as a conseqence of an earlier change which he made so that the OB would be more robust to errors (communication errors, e.g.). Before, when there was an error in presetting, exposing or reading out, the image would not be saved and the OB would crash. Now, when an error is encounter and the exposure is lost, it is flagged as such, and the OB continues to the end but then starts again, skipping all good dither positions and expsoures until it reaches the lost one which it then repeats.

afternoon DX is working now. Aaron & Butch went into the cell. Aaron said that the board for one of the HP is suspect.

conditions Sky is mostly clear but there is some very sparse cirrus - not strictly photometric.

17:15 both LBCs filled

17:25 Tim got the temperature and humidity sensor back up and running.

17:30 Logged in as LBTO, copied OBs to /home/LBTO/0ObservingBlocks/ (leading "0" to make the directory first in the list)

17:30 turning on LBC and taking a few biases while Aaron is opening up the enclosure. B,R are 003849,003847 and slewing to zenith position (so 004... may have banding). Approximate bias levels:
LBC-Blue: 1: 770 2: 850 3: 770 4: 750
LBC-Red: 1: 420 2: 560 3: 370 4: 360

17:50 1-sec I-BESSEL flats (SkyFlat_BI_5) 005118,005118 ---argh red side only used chip 2 (error in OB). Repeat at 5-sec.

18:00 1-sec UY test. It is really too dark for these filters but try with 60sec. 8000->4000 cnts in R.

18:13 60sec r (B) and 30sec r (R) flats. OB was set for 1sec and 0.5 set at B and R, respectively.

18:24 Slewing to SA 92 to collimate & obtain calibration data. not strictly photometric - field used for other purposes, and astrometric fields not yet high

18:30 Clearing coefficients & setting Z11 (SX) = +1000.

18:41 Finished collimating (5 iterations and manual kick to R initially, for Z4=10,000). Polaris seeing 1.04" but LBCFPIA estimates are 1.58(B) and 1.02(R). Taking SA92_UBVVRI to check image quality and stability. 014833 is elongated (ell ~ 0.5) but corresponding blue (014813) is round (fwhm~5pxl = 1.1")

18:50 Check rotator center now that DX is collimating well. Rotator center is just 1pix x,y off from what I measured last night.


      Old rotator center: 1079. 2893.
      New rotator center:
      ##################################################
      Xc=1072.78 Yc=2916.44 from geomap.
      Xc=1072.65 Yc=2916.57 from averages.
      ##################################################
      Images are 2304 pixels wide

used first PA=0 image and got (1073, 2917) while second PA=0 image gave (1073, 2916)

19:30 Taking tilt data. RB_rV superfoc PA=0 023747...024033, PA=180 024413...024656

some gradient on PA=0 chip2 but not on PA=180 chip 2 needs more analysis

20:18 Near zenith at elev=87, hit X=-2.6 limit on DX. Mirror panicked. Raised mirror. Slewed down to lower elevation and am collimating here.

21:00 We entered Global Offsets of X=1.2mm, RY=-25.24" to get mirror out of -ve X limit at high (87 deg) elevation. This has changed the pointing and it is clear that we are not co-pointed. Going to a pointing star. Do we need to make the pointing model for R with the global offsets. DX already had Y, RX global offsets and now it has X,RY ones.

(The log was completed in a txt file and mailed around).

-- OlgaKuhn - 22 Dec 2008
Topic revision: r3 - 21 Jul 2009, NormCushing
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback